# Benchmarking Computer Science Survey Generation

### v1

@@ 4.3 Structural Quality
A clear and logical structure is fundamental to a high-quality scientific survey, as it greatly enhances readability, coherence, and the ability to quickly navigate information. In contrast, poor structural organization can cause confusion, redundancy, and difficulty in following the paper’s arguments. Therefore, a robust evaluation of structural quality is essential. We propose to assess structure from two complementary perspectives: coarse-grained, high-level organization and fine-grained, heading-level alignment.
To this end, we introduce two metrics for structural assessment:
  • Structure Quality Score (SQS): Assesses the high-level structural alignment by comparing the list of generated headings with the ground-truth (GT) headings, focusing on their similarity in structure, meaning, and wording.
  • Soft-Heading Recall (SHR): Measures the fine-grained alignment of headings and subheadings by using semantic embeddings to account for partial or paraphrased matches.
@@
SQS is designed to evaluate the overall quality of the generated survey’s outline. For this evaluation, we define an outline as a hierarchical list of section and subsection headings extracted from a survey. SQS assesses the high-level structural alignment by comparing the generated outline against the ground-truth (GT) outline, focusing on aspects like topical coverage, logical flow, and hierarchical organization. To compute this score, we employ the "LLM-as-a-Judge" paradigm. [^3] The corresponding prompt templates are available in our official GitHub repository: https://github.com/oneal2000/SurGE. Specifically, we prompt a powerful large language model, GPT-4o, with both the generated and the GT outlines. The model is then instructed to provide a holistic quality score from 1 (poor) to 5 (excellent).
@@
To measure fine-grained alignment, SHR evaluates how well the generated outline covers the specific headings present in the ground-truth outline. Unlike metrics based on exact lexical matching, SHR leverages semantic similarity to robustly handle variations in wording and paraphrasing.
Formally, SHR is defined as the soft cardinality overlap between the predicted heading set ($H_{P}$) and the ground-truth heading set ($H_{GT}$):
SHR $=\frac{\mathcal{S}(H_{P}\cap H_{GT})}{\mathcal{S}(H_{GT})}$ (3)
where $\mathcal{S}(A)$ denotes the soft cardinality of a heading set $A$. It is designed to down-weight redundant headings within the same set and is calculated as:
$\mathcal{S}(A)$ $=\sum_{i=1}^{K}\frac{1}{\sum_{j=1}^{K}\text{sim}(A_{i},A_{j})}$ (4)
Here, $\text{sim}(A_{i},A_{j})$ is the cosine similarity between the embeddings of headings $A_{i}$ and $A_{j}$:
$\text{sim}(A_{i},A_{j})$ $=\cos(\text{embed}(A_{i}),\text{embed}(A_{j}))$ (5)
Since a standard set intersection would be too strict for comparing paraphrased headings, we calculate the soft cardinality of the intersection $\mathcal{S}(H_{P}\cap H_{GT})$ using the soft set inclusion-exclusion principle:
$\mathcal{S}(H_{P}\cap H_{GT})=\mathcal{S}(H_{P})+\mathcal{S}(H_{GT})-\mathcal{S}(H_{P}\cup H_{GT})$ (6)
This formulation ensures that semantically matched headings contribute positively to the recall, even with lexical variations. A higher SHR score indicates better structural alignment with the GT outline at a granular level.
@@ 4.4 Content Quality
Content quality evaluation is pivotal for ensuring surveys are not only comprehensive but also clear, coherent, and logically sound. Evaluating this dimension involves assessing the presence of essential information, accuracy, and fluency of expression. While traditional metrics such as ROUGE and BLEU effectively measure textual overlap and fidelity, they inadequately capture nuanced aspects of logical coherence and clarity. To address these limitations, we supplement these traditional metrics with a comprehensive logical coherence evaluation.
Metric Definitions.
We apply ROUGE and BLEU at the section-level to effectively capture fidelity and precision in content representation without needing explicit mathematical formulations here. Logical coherence is evaluated using GPT-4o to assign readability and coherence scores (0–5) for each section.  [^4] Here, we employ the LLM-as-a-Judge paradigm for evaluation. The corresponding prompt templates are available on our official GitHub repository: https://github.com/oneal2000/SurGE.
5 Baseline Implementation
In this section, we detail the implementation of baseline methods. Each baseline follows a two-stage pipeline: (1) retrieving a set of potentially relevant papers for a given topic, and (2) organizing and summarizing the retrieved content to produce a structured survey. To ensure a fair comparison, all baselines rely on a shared dense retriever for the first stage. In the following subsections, we first describe the training procedure of the shared Paper Retriever, followed by the specific methodologies adopted by each baseline in generating the final survey.
@@ 5.1 Paper Retriever Training
We employ a dual-encoder architecture for document retrieval, initialized with roberta-base. For a query $q$ and a paper abstract $d$, we first construct their input representations by prepending the special token $\mathrm{CLS}]$ and appending $\mathrm{SEP}]$. Formally, let $X(q)=[\mathrm{CLS}]\,q\,[\mathrm{SEP}]$ and $X(d)=[\mathrm{CLS}]\,d\,[\mathrm{SEP}]$. We then feed these tokens into the transformer encoder to obtain a hidden vector at the $\mathrm{CLS}]$ position:
@@
The similarity score between the query and the document is computed as the dot product of their embeddings:
@@
During training, each query $Q$ is paired with a relevant document $d^{+}$ to form a positive sample, while negative samples $d^{-}\in N$ are drawn from the corpus. The retriever is optimized via the softmax cross-entropy loss:
@@
This objective encourages the model to assign higher scores to relevant papers while minimizing scores for irrelevant ones. We use the trained retriever to retrieve the top-ranked papers for each query, thereby providing a focused subset of literature for subsequent summarization.
5.2 Retrieval-Augmented Generation
5.3 AutoSurvey
AutoSurvey wang2024autosurvey implements a multi-stage survey generation pipeline that starts with a high-level outline and proceeds through iterative expansions. We adapt it to use our fine-tuned retriever in place of its original retrieval mechanism and keep the number of retrieved references consistent for fairness. In the adapted workflow, we first issue a query based on the topic $t$ to retrieve an initial collection of papers $P_{\text{init}}$. The LLM is then prompted to create a structured outline, which includes main sections and subsections tailored to the subject matter. Next, each section is expanded by conditioning on the subset of papers most relevant to that specific section, producing a draft that includes references in a bracketed format (e.g., “[id]”). Once each section is drafted, the LLM refines it to address factual inconsistencies, stylistic mismatches, and reference-formatting issues. Finally, all refined sections are merged into a coherent final survey, with transitions and citation references carefully aligned. The workflow iterates over these stages, leading to incremental improvements in both thematic coverage and presentation quality.
5.4 StepSurvey
StepSurvey lai2024instruct is a more granular generation strategy that also begins with retrieving the top 100 candidate papers for a given topic $t$ but proceeds through distinct planning and drafting phases in a sequential manner. This baseline is proposed by a team named “ID” in the NLPCC2024 competition task 6 tian2024overview . Rather than producing an overarching outline at once, it starts by proposing a survey title and a set of primary headings that collectively capture the central themes of the retrieved literature. Subsequently, it uses the primary headings to guide the selection of secondary or finer-grained topics, each mapped to a relevant subset of the retrieved papers. The LLM then produces a full draft by writing each subsection with explicit attention to references and academic conventions, thereby encouraging greater control and consistency across sections. Throughout this process, the system attempts to maintain a balanced level of detail, striving for a clear exposition of important subtopics while avoiding excessive verbosity or redundancy. By structuring the content in incremental steps, StepSurvey aims to achieve coherent organization and thorough coverage of the literature.
@@ 5.4 StepSurvey
For the training of Paper Retriever, we randomly split the dataset into a training set and a test set at a ratio of 4:1. We adopt the AdamW optimizer for model optimization, the learning rate is set to $5\times 10^{-6}$, the weight decay is set to $1\times 10^{-2}$, and the epoch is set to $10$. During the training process, we adopt mixed-precision (fp16) training for enhanced efficiency. At inference time, each query retrieves the top 100 most relevant papers according to the similarity score $S(q,d)$. The retriever is initialized using the pre-trained RoBERTa model liu2019roberta . For the generation configuration of LLMs, all experiments are conducted using the publicly available implementations provided by Hugging Face. We utilize the default hyperparameters and the chat template as outlined in the official Hugging Face repository  [^5] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct. All experiments are conducted on a server with 1TB of RAM and eight NVIDIA A100 GPUs, each with 40GB of memory.
@@ 6 Experimental Results
This section demonstrates the utility of our proposed benchmark by presenting a detailed evaluation of three representative survey generation baselines: RAG, AutoSurvey, and StepSurvey. Our analysis is designed to showcase how the benchmark’s multi-faceted metrics can be used to conduct a fine-grained analysis and pinpoint system bottlenecks. We first isolate and evaluate the performance of the shared Paper Retriever (§6.1). This analysis is crucial as it establishes the theoretical upper-bound performance for the reference coverage metric and highlights the challenges posed by the initial retrieval stage. Following this, we assess the complete end-to-end performance of the baselines (§6.2), leveraging our comprehensive evaluation suite to compare their architectural strengths and weaknesses and to shed light on the broader challenges facing automatic survey generation.
@@ 6.1 Analysis of Retrieval Performance
A crucial question in our two-stage pipeline is whether performance limitations stem from the retriever’s inability to find relevant papers or the generator’s inability to use them. To disentangle these factors and quantify the retrieval bottleneck, we first evaluate the performance of our fine-tuned dense retriever in isolation. This analysis establishes the theoretical upper bound for the reference coverage that our end-to-end systems can achieve.
We compare our dense retriever against a lexical baseline BM25 robertson2009probabilistic , using Recall@k as the evaluation metric. This metric measures the percentage of ground-truth papers from the reference survey that are present in the top-k retrieved documents. As shown in Table 1, our fine-tuned Paper Retriever substantially outperforms the BM25 baseline across all values of k. The performance gap underscores the inadequacy of lexical search, which struggles to find semantically relevant papers that may not share overlapping keywords. Our dense retriever, by capturing deeper semantic relationships, is far more effective. However, the results also reveal a critical bottleneck in the survey generation pipeline. Even with a generous budget of k=1000, Paper Retriever recalls only 68.05% of the ground-truth papers. This finding is crucial as it establishes a hard upper-bound on the performance of the downstream generator. This finding suggests that achieving a breakthrough in survey quality may necessitate a shift towards more sophisticated retrieval paradigms, such as employing search agents powered by large language models zhang2024large .
Table 2: The experimental results combining Coverage, Referencing Accuracy (including Doc-Rel, Sec-Rel, and Sent-Rel), Structural Quality (including SQS and SHR), and Content Quality (as measured by ROUGE scores, BLEU, and Logic Score). Doc-Rel represents document-level relevance, Sec-Rel represents section-level relevance, Sent-Rel represents sentence-level relevance, SQS represents the structure quality score, and SHR represents soft-heading recall. R-L represents ROUGE-L. The best results are in bold, and the second-best results are underlined.
Coverage Referencing Accuracy Structural Quality Content Quality
Baseline Recall Doc-Rel Sec-Rel Sent-Rel SQS SHR R-L BLEU Logic Score
@@ 6.2 Overall Performance of Survey Generation Baselines
The preceding analysis revealed that the retriever provides the generation models with over a third (36.65%) of the ground-truth references, setting a clear performance ceiling. We now demonstrate the full diagnostic power of our benchmark by applying its comprehensive suite of metrics to the end-to-end outputs of the three baselines: RAG, AutoSurvey, and StepSurvey. As detailed in Table 2, this multi-faceted evaluation allows us to pinpoint specific challenges in the generation stage. The key insights revealed by our benchmark are as follows:
(1) Bottleneck of Existing Baselines. A striking observation across all methods is the significant gap between retrieval performance and final survey coverage. While our paper retriever makes 36.65% of the ground-truth references available (Recall@100), the best-performing baseline, StepSurvey, only achieves a final Coverage of 6.30%. The simpler RAG baseline is even lower at 2.14%. This disparity underscores that the primary limitation is the generation stage’s inability to effectively identify and incorporate relevant information from the provided set of 100 papers. Nonetheless, the results also show that advanced, structured approaches like AutoSurvey (3.51%) and StepSurvey (6.30%) are substantially better at utilizing the retrieved context than a standard RAG pipeline.
(2) Strengths of AutoSurvey. AutoSurvey exhibits excellent Section-Level Relevance (0.4935) and Sentence-Level Relevance (0.4870), surpassing both RAG and StepSurvey. This improvement appears to stem from AutoSurvey’s iterative planning mechanism, where the system generates a high-level outline, retrieves relevant references for each section, and incrementally refines each subsection. By focusing on context-specific references and refining drafts section by section, AutoSurvey places citations more accurately at both the section and sentence levels. Moreover, AutoSurvey achieves the highest Structure Quality Score (SQS) of 1.3902, indicating that its top-down planning leads to well-organized hierarchical structures.
(3) Strengths of StepSurvey. On the other hand, StepSurvey attains the best performance in several crucial dimensions: Coverage (6.30%) and Document-Level Relevance (0.4576). Its multi-phase workflow first generates major headings, then subtopics, and finally assigns references to each subsection. This step-by-step refinement appears particularly effective at broadening the coverage of included works and aligning the cited literature with the overarching survey theme. StepSurvey also achieves the highest Logic Score (4.8451) and excels in ROUGE and BLEU, reflecting strong content quality and coherent presentation. An interesting observation is that while StepSurvey’s Section-Level Relevance is slightly lower than AutoSurvey’s, the difference remains small (0.4571 vs. 0.4935), indicating that its primary strength lies in more holistic, document-wide alignment and consistent logical flow.
(4) Impact of Planning on Multi-Level Relevance. The multi-stage generation process adopted by AutoSurvey and StepSurvey benefits Document-Level and Section-Level Relevance in different ways. AutoSurvey’s approach—outlining, drafting, and iteratively refining sections—yields highly relevant citations at the subsection and sentence granularity. In contrast, StepSurvey’s method of organizing subtopics around central themes ensures strong document-wide alignment. This discrepancy suggests that carefully balancing local vs. global planning remains an open challenge: future work might explore hybrid strategies that combine AutoSurvey’s fine-grained focus with StepSurvey’s broad thematic coverage.
In conclusion, our results demonstrate that advanced, multi-stage planning approaches (AutoSurvey and StepSurvey) significantly outperform the standard RAG pipeline in nearly all metrics. While AutoSurvey excels in section- and sentence-level citation placement and high-level structural quality, StepSurvey provides superior document-level coverage, global relevance, and strong logical coherence. Nonetheless, all methods still struggle to capture the vast reference space fully. These findings underscore the importance of iterative refinement and structured planning in automated survey generation and suggest that there remains substantial room for further improvement, particularly in boosting coverage and refining the interplay between local and global survey organizations.

### v2

@@ 4.3 Structural Quality
The logical flow and coherence of a scientific survey are fundamentally determined by its structure. This makes structural quality a critical factor for readability and overall impact. To comprehensively evaluate structural quality, we introduce two complementary metrics that assess the generated outline at both macroscopic and microscopic levels. Our first metric, the Structure Quality Score (SQS), addresses the high-level organization. It holistically assesses the alignment between the generated and ground-truth outlines by comparing their overall structure, semantic coherence, and topical progression. Complementing this, our second metric, Soft-Heading Recall (SHR), provides a fine-grained evaluation of heading alignment. It specifically measures how well the generated headings cover those in the ground-truth based on the similarity of semantic embeddings.
@@
SQS evaluates the overall quality of a generated survey’s structure based on the hierarchical list of its section headings. To compute this score, we adopt the LLM-as-a-Judge paradigm, leveraging GPT-4o as the evaluator. Specifically, we provide the LLM with both the generated and ground-truth outlines and prompt it to assign a quality score. To guide the LLM’s evaluation, we have carefully designed a detailed instruction prompt that includes a comprehensive scoring rubric on a scale from 0 to 5. The complete prompt is shown in Appendix E.
@@
To measure fine-grained alignment, SHR evaluates how well the generated outline covers the specific headings present in the ground-truth outline. Unlike metrics based on exact lexical matching, SHR leverages semantic similarity to robustly handle variations in wording and paraphrasing. Formally, SHR is defined as the soft cardinality overlap between the predicted heading set ($H_{P}$) and the ground-truth heading set ($H_{GT}$):
SHR $=\frac{\mathcal{S}(H_{P}\cap H_{GT})}{\mathcal{S}(H_{GT})},$ (4)
where $\mathcal{S}(A)$ denotes the “soft cardinality” of a heading set $A$. Intuitively, this metric counts the number of semantically unique headings in a set. It achieves this by down-weighting redundant headings. Specifically, the contribution of each heading is inversely proportional to its aggregated similarity with all other headings in the set:
$\mathcal{S}(A)$ $=\sum_{i=1}^{K}\frac{1}{\sum_{j=1}^{K}\text{sim}(A_{i},A_{j})}.$ (5)
Here, $\text{sim}(A_{i},A_{j})$ is the cosine similarity between the embeddings of headings $A_{i}$ and $A_{j}$. A standard set intersection would be too strict for comparing paraphrased headings. Therefore, we define the soft intersection cardinality using the inclusion-exclusion principle:
$\mathcal{S}(H_{P}\cap H_{GT})=\mathcal{S}(H_{P})+\mathcal{S}(H_{GT})-\mathcal{S}(H_{P}\cup H_{GT}).$ (6)
The core idea lies in the union term, $\mathcal{S}(H_{P}\cup H_{GT}$. When computed on the combined heading set, a predicted heading and a similar ground-truth heading mutually reduce the union’s soft cardinality. This reduction directly quantifies their semantic overlap, allowing the metric to reward paraphrased matches. A higher SHR score thus indicates better granular alignment.
@@ 4.4 Content Quality
To assess the content quality of generated scientific surveys, we propose the Content Quality Score (CQS) metric based on the LLM-as-a-Judge paradigm, leveraging GPT-4o to evaluate each section of the survey. The evaluation is based on five criteria: fluency and coherence, logical clarity, avoidance of redundancy, clarity of description, and absence of errors. To guide the LLM’s evaluation, we designed a detailed instruction prompt for the LLM, which is provided in the Appendix E. Each section is scored on a scale of 0 to 5, where a higher score reflects superior fluency, logical progression, and clarity. Considering the context length limitations of the LLM, we have it score each survey section by section, and the final score is the average of the scores from all sections.
As supplementary measures, we also compute ROUGE and BLEU scores, which quantify n-gram overlap between the generated and ground-truth surveys. While these are well-established metrics in text generation, their role is secondary in our framework, serving as additional checks for content fidelity rather than a primary assessment method.
5 Experimental Setup
In this section, we detail the implementation of our experiment. Each baseline follows a two-stage pipeline: (1) retrieving a set of potentially relevant papers for a given topic, and (2) organizing and summarizing the retrieved papers to produce a structured survey. For fair comparison, all baselines share the same dense retriever for the first stage. In the following subsections, we first describe the training of the shared Paper Retriever, followed by the baseline selection and implementation details.
@@ 5.1 Paper Retriever Training
We employ a dual-encoder architecture for retrieval, initialized with roberta-base. The training process leverages the benchmark dataset, where each topic description $t$ (introduced in §2) serves as the query $q$. For a query $q$ and a paper abstract $d$, we construct their input representations by prepending the special token $\mathrm{CLS}]$ and appending $\mathrm{SEP}]$. Formally, let $X(q)=[\mathrm{CLS}]\,q\,[\mathrm{SEP}]$ and $X(d)=[\mathrm{CLS}]\,d\,[\mathrm{SEP}]$. We then feed these tokens into roberta-base to obtain the contextualized embedding of the $\mathrm{CLS}]$ token:
@@
The similarity score between the $q$ and the $d$ is computed as the dot product of their embeddings:
@@
During training, each query $Q$ is paired with the relevant documents $d^{+}$ from the ground truth paper to form the positive samples, while negative samples $d^{-}\in N$ are randomly sampled from the corpus. The retriever is optimized via the softmax cross-entropy loss:
@@
This objective encourages the model to assign higher scores to relevant papers while minimizing scores for irrelevant ones. After training, we use the trained retriever to retrieve the top-ranked papers for each query, thereby providing a collection of relevant papers for the subsequent generation stage.
5.2 Baselines
We selected the following three survey generation baseline methods for our experiments. Detailed descriptions of each method are provided in the appendices.
  • Retrieval-Augmented Generation (RAG): This straightforward approach retrieves relevant papers, summarizes them in chunks, and then merges these summaries to form the final survey. A detailed description is provided in Appendix D.1.
  • AutoSurvey (Wang et al., 2024d): This method employs a multi-stage, outline-driven pipeline. It first generates a high-level outline from retrieved papers and then iteratively expands and refines each section. Further details can be found in Appendix D.2.
  • StepSurvey (Lai et al., 2024b): This approach adopts a granular, step-by-step process. It starts by generating a title and primary headings, which then guide the incremental drafting of each subsection. The methodology is described in detail in Appendix D.3.
@@ 5.2 Baselines
For the training of Paper Retriever, we randomly split the dataset into a training set and a test set at a ratio of 4:1. We adopt the AdamW optimizer for model optimization, the learning rate is set to $5\times 10^{-6}$, and the epoch is set to $10$. During the training process, we adopt mixed-precision (fp16) training. At inference time, each query retrieves the top 100 relevant papers according to the similarity score. The retriever is initialized using the pre-trained RoBERTa model (Liu et al., 2019).
To ensure a fair comparison, we utilize the Qwen2.5-14B-Instruct model (Yang et al., 2024a) as the base LLM for all baseline methods. For the generation configuration of LLMs, all experiments are conducted using the publicly available implementations provided by Hugging Face. We utilize the default hyperparameters and the chat template as outlined in the official Hugging Face repository  [^4] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct. All experiments are conducted on a GPU server with 1TB RAM and eight NVIDIA A100 GPUs, each with 40GB of memory.
@@ 6 Experimental Results
This section presents the detailed evaluation of the selected survey generation baselines. We first evaluate the performance of the shared Paper Retriever (§6.1). This analysis is crucial as it establishes the theoretical upper-bound performance for the comprehensiveness metric and highlights the challenges posed by the initial retrieval stage. Following this, we assess the complete end-to-end performance of the three baselines (§6.2) based on our proposed evaluation framework.
@@ 6.1 Analysis of Retrieval Performance
A crucial question in our two-stage pipeline is whether performance limitations stem from the retriever’s inability to find relevant papers or the generator’s inability to use them. To disentangle these factors and quantify the retrieval bottleneck, we evaluate the performance of our fine-tuned dense retriever in isolation. This analysis establishes the theoretical upper bound for the reference coverage that our end-to-end systems can achieve.
We compare our dense retriever against the lexical baseline BM25 (Robertson et al., 2009), using Recall@k as the evaluation metric. This metric measures the percentage of ground-truth papers from the reference survey that are present in the top-k retrieved documents. As shown in Table 1, our fine-tuned Paper Retriever substantially outperforms the BM25 baseline across all values of $k$. The performance gap underscores the inadequacy of lexical search, which struggles to find semantically relevant papers that may not share overlapping keywords. On the other hand, our dense retriever can capture deeper semantic relationships, thus it is far more effective than BM25. However, the results also reveal a critical bottleneck in the survey generation pipeline. Even when retrieving the top 1000 documents (k=1000), Paper Retriever’s recall reaches only 68.05% of the ground-truth papers. This performance ceiling imposes a hard upper bound on the downstream generator, as it cannot synthesize information from papers it never receives. Consequently, this performance ceiling underscores that the survey generation task needs more sophisticated retrieval paradigms, such as employing search agents or multi-agent systems powered by large language models (Zhang et al., 2024; Tang et al., 2025; Li et al., 2025).
Table 2: Main experimental results comparing different survey generation baselines across four dimensions: Comprehensiveness (Comp.), Citation Accuracy, Structural Quality, and Content Quality. Metrics include Recall, Document/Section/Sentence-level Citation Accuracy (Doc-Acc, Sec-Acc, Sent-Acc), Structure Quality Score (SQS), Soft-Heading Recall (SHR), ROUGE-L, BLEU, and Content Quality Score (CQS). The best results are in bold, and the second-best results are underlined.
Comp. Citation Accuracy Structural Quality Content Quality
Baseline Recall Doc-Acc Sec-Acc Sent-Acc SQS SHR R-L BLEU CQS
@@ 6.2 Overall Performance of Survey Generation Baselines
Our evaluation now transitions from retrieval performance to the end-to-end survey generation task. While the retriever supplies the models with a substantial portion (36.65%) of the ground-truth references, a detailed analysis of the final surveys reveals a significant bottleneck in the generation stage. Table 2 presents a multi-faceted comparison of our three baselines across the four dimensions of survey quality defined by our benchmark. This comprehensive evaluation yields several key insights:
(1) The Generation Stage is the Primary Bottleneck. A striking finding is the dramatic performance drop-off between paper retrieval and final survey synthesis. Despite the retriever providing access to 36.65% of ground-truth references (Recall@100), the best-performing model, StepSurvey, incorporates only 6.30% of them into its final output. The standard RAG baseline fares even worse, with a final recall of just 2.14%. This stark disparity highlights that the core challenge lies not in finding relevant papers, but in the generator’s ability to effectively identify and integrate crucial information from the retrieved set. Nevertheless, the results also affirm the value of structured pipelines, as AutoSurvey (3.51%) and StepSurvey (6.30%) significantly outperform the naive RAG approach in comprehensiveness.
(2) AutoSurvey Excels in Local Coherence and Structural Quality. AutoSurvey demonstrates particular strength in producing locally coherent and well-organized content. As shown in Table 2, it achieves the highest scores for both Section-level Citation Accuracy (Sec-Acc: 0.4935) and Sentence-level Citation Accuracy (Sent-Acc: 0.4870). This suggests its iterative, section-by-section refinement process is highly effective at placing citations within their correct immediate context. Furthermore, their planning approach results in the highest Structure Quality Score (SQS) of 1.3902, indicating superior hierarchical organization in the generated survey.
(3) StepSurvey Achieves Superior Coverage and Global Content Quality. In contrast, StepSurvey’s strengths lie in metrics related to holistic coverage and overall content quality. It attains the highest Recall (0.0630) and Document-level Citation Accuracy (Doc-Acc: 0.4576), demonstrating a better ability to cover the breadth of the topic and align cited works with the survey’s main theme. Its multi-phase workflow that generates headings, then subtopics, and finally assigns citations leads to the best content quality. This is evidenced by its top scores in text similarity metrics (ROUGE-L: 0.1590, BLEU: 12.02) as well as the highest Content Quality Score (CQS) of 4.8451, which reflects stronger logical flow and presentation.
(4) A Trade-off Between Local Precision and Global Coverage. The distinct strengths of AutoSurvey and StepSurvey reveal a fundamental trade-off in planning strategies. AutoSurvey’s iterative refinement of individual sections fosters high-quality local structure and precise, fine-grained citation placement. Conversely, StepSurvey’s hierarchical, topic-first approach yields better overall topic coverage and more coherent, globally relevant content. While both advanced methods significantly outperform the standard RAG pipeline, neither excels across all dimensions.
In conclusion, our proposed SurGE benchmark effectively diagnoses the weaknesses of current survey generation systems. The results clearly indicate that advanced, multi-stage planning is crucial, but a substantial gap remains between machine-generated and expert-written surveys, particularly in comprehensiveness. The observed trade-off between local and global optimization strategies highlights a key challenge for future research: developing hybrid models that can combine the fine-grained accuracy of AutoSurvey with the broad thematic coverage and coherence of StepSurvey.